### v1

# Enhancing the Efficiency of Complex Systems Crystal Structure Prediction by Active Learning Guided Machine Learning Potential

Abstract
Understanding multicomponent complex material systems is essential for design of advanced materials for a wide range of technological applications. While state-of-the-art crystal structure prediction (CSP) methods effectively identify new structures and assess phase stability, they face fundamental limitations when applied to complex systems. This challenge stems from the combinatorial explosion of atomic configurations and the vast stoichiometric space, both of which contribute to computational demands that rapidly exceed practical feasibility. In this work, we propose a flexible and automated workflow to build a highly generalizable and data-efficient machine learning potential (MLP), effectively unlocking the full potential of CSP algorithms. The workflow is validated on both Mg–Ca–H ternary and Be–P–N–O quaternary systems, demonstrating substantial machine learning acceleration in high-throughput structural optimization and enabling the efficient identification of promising compounds. These results underscore the effectiveness of our approach in exploring complex material systems and accelerating the discovery of new multicomponent materials.
I Introduction
Functional materials with unique physical and chemical properties are essential for advancing cutting-edge technologies and driving breakthroughs across various disciplines[1, 2, 3, 4, 5]. Compared to simple systems, multicomponent materials with intricate atomic configurations and vast compositional diversity enable fine-tuning of electronic, magnetic, and structural properties, thus offering significantly greater opportunities for discovering a wide range of high-performance functional materials[6, 7]. Examples include high superconductors[8, 9, 10, 11], multifunctional alloys[12, 13], high ionic conductivity solid electrolytes[14], and nonlinear optical materials[15, 16, 17, 18]. However, the design of multicomponent complex materials remains a major challenge. The combinatorial explosion results in highly complex thermodynamic phase diagrams, while structural diversity[19] and the scarcity of comprehensive data [20] further hinder both experimental and computational exploration, particularly when materials are expected to exhibit both thermodynamic synthesizability and superior functional properties. Given the vast combinatorial complexity of arranging chemical elements from the periodic table into all possible stoichiometries, current knowledge likely captures only a small fraction of the immense material design landscape[21]. Therefore, advanced techniques are essential for efficiently navigating this space and fully unlocking the potential of complex material systems.
Among theoretical materials design approaches, crystal structure prediction (CSP)[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] has emerged as an indispensable tool for predicting stable crystalline configurations prior to experimental synthesis[33, 34, 35, 36, 37, 38, 39]. Fundamentally, CSP approaches can be formulated as global (or multi-local) optimization problems on high-dimensional energy surfaces, thereby making accurate free energy calculations essential for reliable predictions. In practice, most conventional CSP implementations employ density functional theory (DFT)[40] as their computational backbone. Despite its reliability, DFT involves solving the electronic structure through an iterative self-consistent field (SCF) procedure, with computational costs scaling cubically with system size ( for electrons). Such unfavorable scaling renders structural searches computationally expensive, thereby restricting most applications to small unit cells with limited compositional complexity. Consequently, the full potential of CSP methods are significantly constrained.
Trained on datasets derived from DFT calculations, machine learning potentials (MLPs) can faithfully reproduce DFT-level potential energy surfaces (PESs) while delivering computational speedups by several orders of magnitude. Research efforts such as GAP-RSS[41], DP-CALYPSO[42], MTP-USPEX[43], and MAGUS[29] demonstrate the feasibility of this paradigm. These approaches typically begin with a concurrent learning sampling scheme to construct a representative DFT dataset along with a corresponding potential model, which is then used to drive CSP for discovering further unknown structures. However, due to the interpolative nature of MLPs, their performance heavily relies on the quality of the dataset. When applied to more complex systems, where the number of relevant local minima increases exponentially[23], assessing whether an MLP can generalize well enough to reliably predict unforeseen configurations is often nontrivial. The risks includes leading CSP to converge to unphysical local minima, thereby compromising the reliability of structure prediction and potentially overlooking meaningful configurations. Recently, an alternative direction has focused on constructing foundation models trained on extensive datasets spanning the whole periodic table[44, 45, 46, 47, 48, 49, 50, 51]. Although significant improvements in generalization capability have been observed, the fundamental limitation remains unresolved, especially with data scarcity under extreme pressure conditions, strongly correlated systems, or low-symmetry and defect-rich systems. A robust approach for rapid and exhaustive structural prediction of multicomponent complex systems remains urgently needed.
In this work, we first implemented an MLP, named the attention coupled neural network (ACNN), which demonstrated both high accuracy and computational efficiency in surrogating the PES. Base on ACNN, we developed an automated pipeline specifically tailored to address the challenges of predicting structures for materials with multicomponent variable stoichiometry. At the core of this pipeline is a flexible active learning process that orchestrates the dynamic training of the ACNN and the CSP procedure, ensuring that the ACNN model consistently maintains strong generalization capability throughout. Concurrently, the continuous identification of new structures drives the ongoing refinement of phase diagrams. We validated our algorithm by exploring the high-pressure, defect-rich Mg–Ca–H ternary and Be–P–N–O quaternary systems across expanded compositional and structural spaces. Our approach enabled the optimization of structures with quantities several orders of magnitude greater than those achievable by DFT, significantly broadening the structural search space. Both cases uncovered a wide range of novel, thermodynamically synthesizable materials, including high-temperature superconductors and nonlinear optical materials, highlighting the high data efficiency and low computational cost of our method.
II Methodology
II.1 Machine learning potential
ACNN incorporates various features extensively discussed in previous studies[52, 53, 54, 55, 56, 57, 58, 59]. In particular, it explicitly describes intrinsic translational, rotational, and permutational invariance for scalar quantities (energy), and rotational equivariance for tensorial properties (atomic forces and stress tensors). Additionally, the consistently continuous and smooth differentiability should also be satisfied. Following Behler and Parrinello [52], the total potential energy is expressed as a sum of atomic contributions .
where denotes the total number of atoms. The force and virial tensors are directly computed from via
The central term is uniquely defined by the local environment of atom ,
in which denotes the relative position vector from atom to atom . ACNN utilizes an analytical descriptor of , whose form is derived from the atomic cluster expansion (ACE)[60, 58] framework. Specifically, the atomic energy is expanded in terms of -body correlation functions,
The energy is obtained through the nonlinear combination effects of the neural network, expressed as
where are model parameters associated with the neural network architecture. Typically, the interaction is limited to a predefined cutoff sphere with radius . To smooth the behavior of atoms at the boundary, an envelope function is defined as
For a central atom , the first-order term, which accounts for pairwise interactions, is given by
where is a function of the distance, defined as
with denoteing the Chebyshev polynomial of the first kind of order . The second-order term, which incorporates angular dependencies, is given by
where is the Legendre polynomial of order , and is the angle between the vectors and . In general, incorporating higher-order terms improves the completeness of the local environment description and thereby enhances accuracy. However, achieving convergence becomes progressively slower, as evidenced by the tests provided in Drautz [60]. To balance computational efficiency and precision, ACNN explicitly incorporates up to 3-body interactions, while delegating the remaining residual errors to the nonlinear neural network for further correction.
The overall ACNN architecture consists of three main components: an embedding module, an optional stack of attention modules, and a fitting module, as illustrated in FIG. 1. The data flows through these components sequentially. The vectorized geometry of the input structures is first processed by the embedding module, which incorporates elemental information to enrich the feature space. Specifically, the embedding module employs two separate multilayer perceptron networks (FIG. 1(b) and (c)) to independently embed and as follows:
After embedding, the resulting features are either directly fed into the fitting module or passed through one or more attention modules before reaching the fitting stage.
The core of each attention module is a multi-head self-attention layer. From a physical perspective, it can be viewed as a weighted reconstruction of the local atomic neighborhood, where the model learns to infer the relative contributions of neighboring atoms. We adopt the standard formulation[61] for constructing the query, key, and value matrices, given by
However, we employ a slightly modified score function,
where is a trainable scalar constrained to be greater than 0.01, introduced to automatically scale the attention scores and yield smoother weighting. The fitting module, composed of a multilayer perceptron, simply receives the processed features and pools them to predict atomic energies, as defined by the formulation in Eqs. 11 and 12.
In summary, the ACNN architecture offers a flexible and straightforward framework for implementing and switching between two distinct MLP designs, both of which satisfy the accuracy requirements for CSP. The primary distinction lies in the trade-off between computational cost and accuracy. Introducing attention modules increases model complexity and the number of trainable parameters, thereby enhancing the learning capacity of ACNN. Furthermore, stacking multiple attention modules implicitly introduces an aggregative effect analogous to that of graph neural networks (GNNs) [62], which extends the effective interaction range among atoms. However, this comes at the cost of higher computational overhead and fundamentally limits the applicability of domain decomposition algorithm for efficient parallelization. Therefore, the attention-based ACNN architecture does not meet the computational requirements for large-scale, multicomponent crystal structure prediction tasks of the complex systems considered in this study. Detailed accuracy and computational speed tests are provided in Section S1 of the Supplementary Materials.
II.2 Training and refining
All parameters in Eq. 5 are trainable through the backpropagation process. Considering that the training set may contain structures with varying numbers of atoms, we define the loss function as follows,
Here, is the minibatch size, and the number of atoms in structure . Hatted quantities (e.g., ) represent reference values, whereas unhatted ones correspond to ACNN predictions. , , and are tunable parameters that flexibly adjust the relative contributions of each component in the loss function. We employed the AdamW[63] optimizer, which incorporates L2 regularization to mitigate overfitting by penalizing large parameter values. This may help improve the generalization performance of the ACNN model when trained on a diverse set of structures broadly sampled from the PES.
When applied to CSP, the training set typically needs to be continuously expanded through the active learning process discussed below. Therefore, ACNN supports restarting training or refining the model from any existing checkpoint with a new or updated dataset. In addition, small perturbations can be introduced to the trainable parameters during refinement to reduce the risk of systematic errors in extrapolated predictions. The underlying intention is for the perturbed model to quickly stabilize in well-sampled regions at the start of refinement, while preserving controlled stochasticity in out-of-distribution regions.
II.3 Active learning
As a data-driven paradigm for interatomic potentials, the performance of ACNN highly depends on the quality of the sampling data. Active learning aims to maintain the predictive accuracy and generalization of the ACNN model during the CSP process, while reducing the number of expensive DFT calculations. According to the principle of maximum information gain, two types of structures are particularly necessary to trigger DFT calculations and should be included in the training set. The first type consists of energy-minimum structures. These samples help reduce the discrepancy between ACNN predictions and DFT reference values, guiding the model toward a more accurate surrogate of the PES. The second stems from unphysical local minima, which typically arise from the model’s insufficient representation of certain regions and therefore require further reinforcement.
The flowchart of the active learning pipeline is illustrated in FIG. 2(a), where two types of data acquisition discussed above are naturally integrated. Upon triggering the workflow, a small set of randomly generated structures undergoes DFT self-consistent field (SCF) calculations, which then form the initial training dataset for ACNN. After training, CSP approach is then carried out on the coarse potential until the number of unphysical structures exceeds a threshold . This indicating that the model’s generalization capability is no longer sufficient for further reliable exploration. The resulting structures are firstly filtered through fingerprint-based[27, 23] deduplication. Subsequently, a thermodynamic phase diagram is constructed based on ACNN accuracy. Despite being incomplete or suboptimal, early ACNN-derived phase diagrams still offer useful guidance for sampling and model refinement. Here, we adopt a composition-wise selection strategy: we first identify the compositions with the lowest convex hull enthalpies (), and then select the lowest-enthalpy structures from each of these compositions. These structures, along with those that fail to converge during optimization, are subsequently subjected to DFT calculations.
The active learning workflow proceeds iteratively, gradually improving the accuracy and generalization of the ACNN model. As its performance advances, more structures are discovered and incorporated into the training set, further reinforcing the model’s capability, as schematically illustrated in FIG. 2(b). Since the sampled dataset also serves as the target for structure prediction, the resulting phase diagram or energy sequence at DFT-level accuracy can be directly constructed from the training data. After sufficient iterations, the phase diagram tends to converge, indicating that the search is approaching exhaustion. Once no additional high value structures are identified, the workflow can be terminated.
Since prediction errors are intrinsic to MLPs, evaluating their influence on the reliability of structure prediction is essential. From the perspective of uncertainty quantification, the error can be decomposed into epistemic and aleatoric components[64]. By systematically sampling underexplored regions of PES, the active learning scheme provides an effective mechanism for reducing epistemic uncertainty, the associated errors can be quantitatively assessed on test sets composed of labeled structures from subsequent iterations. Sources of aleatoric uncertainty mainly contributed by the intrinsic training error of the ACNN approach. Efforts to reduce this uncertainty involve preferentially sampling high-value regions, such as local minima, while avoiding configurations that are far from force equilibrium. Collectively, these errors ultimately manifest as inaccuracies in the predicted energies, forces, and stresses during structure optimization. The energy error, , propagates into the calculation of formation enthalpies, thereby affecting the relative stability of structures within the phase diagram. The force error, , determines the equilibrium configurations obtained via structural optimization and influences the extent of structural distortion. The stress error, , manifests in deviations of lattice parameters and cell geometry, and further impacts the enthalpy through its contribution to the term. If these errors are viewed as perturbations to the equilibrium structures, the resulting phase diagrams tend to overestimate the enthalpy of the structures. To achieve more accurate phase boundaries, it is often necessary to perform full DFT structural relaxations for the structures near the convex hull.
II.4 Computation methods
To assess the effectiveness of the ACNN-based active learning approach in complex material system design, we have conducted systematic tests on two variable composition chemical systems: (i) ternary Mg-Ca-H system, and (ii) quaternary Be-P-N-O system. All known elementary, binary, ternary, and quaternary structures are used as reference structures to evaluate of the thermodynamic stability of newly discovered structures, however, none of them are included a priori in the training set. This ensures a fairer evaluation by more closely reflecting practical application scenarios.
The ACNN hyperparameter configurations for both systems share the following settings: elemental embedding networks with [64, 64] configuration (two layers of 64 neurons), fitting networks with [256, 256] configuration (two layers of 256 neurons), radial basis functions of order , angular basis functions of order , and an angular momentum cutoff of for characterizing single-bond interactions. The cutoff distances differ between systems: for the Mg-Ca-H system under high pressure, Å and Å are used, whereas for the Be-P-N-O system, Å and Å . In each iteration of the active learning workflow, the ACNN is trained for backpropagation steps. An exponential decay learning rate schedule is adopted, with the learning rate decaying by a factor of 0.97 every steps.
To achieve optimal computational efficiency of ACNN, we employ an in-house implementation of the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm to perform batch structure optimization. The batch mode enables multiple structures to be optimized simultaneously, significantly increasing the degree of parallelism, ensuring efficient utilization of computational resources during the CSP workflow. The convergence criteria were set at an energy threshold of eV/atom and a force threshold of eV/Å for each directional component.
For structure generation, we employ the Crystal Structure Analysis by Particle Swarm Optimization (CALYPSO) code[26, 27] in a purely random sampling mode with only symmetry constraints imposed. This ensures an unbiased evaluation of our approach’s performance by eliminating methodology-specific biases inherent to conventional CSP techniques. Though, the swarm-intelligence-based evolutionary approach has been demonstrated to be highly efficient in practical applications.
All DFT calculations were performed using the Ab initio atomic mateRial modEling Software (ARES) package [65], within the framework of the generalized gradient approximation level of theory[40], and employing the optimized norm-conserving Vanderbilt pseudopotentials [66, 67] to account for core-electron effects. For the Mg-Ca-H system, the plane-wave and density cutoff energies were set to 70 Ry and 280 Ry, respectively, while for the Be-P-N-O system, they were set to 80 Ry and 320 Ry. Both systems share the same k-point spacings configuration, with for SCF calculations during structure prediction and for electronic structure calculations.
III Results and discussion
III.1 Mg-Ca-H system
Recent breakthroughs in superconductivity research have been achieved through the theoretical prediction of clathrate-type hydride compounds under high pressure, such as [37, 38, 39] and [68, 34, 36, 35, 69], which were later confirmed by experimental synthesis. These materials, known as “metal superhydrides”, feature a distinctive structural motif where metal atoms are encapsulated within hydrogen-based clathrate cages. As one of the pioneering examples, exhibits a high superconducting critical temperature () of 215 K at 172 GPa. Theoretical frameworks suggest that lighter metal atoms exhibit lower frequency vibrational modes that favor superconductivity, along with higher electron-phonon coupling from weaker multicentered hydrogen bonding interactions[70]. These insights have motivated the incorporation of Mg into the Ca–H binary system, forming the ternary Mg–Ca–H system as a potential platform for high-temperature superconductivity.
Previous study by Sukmas et al. [71] reported a metal-substituted phase, Mg0.5Ca0.5H6, derived from the novel CaH6 prototype. This phase was predicted to exhibit near room-temperature superconductivity, with an estimated of 288 K at 200 GPa. More recently, Cai et al. [72] successfully synthesized Ca-Mg-based ternary superhydrides using binary intermetallic CaMg2 and 1:1 Ca-Mg mixture as starting reactants. The resulting compounds exhibited approaching 168 K at 310 GPa in the CaMg2-based superhydride and 182 K in 1:1 the Ca-Mg superhydride at 324 GPa. These findings indicate the existence of high compound in this system. However, due to the prohibitive computational cost of exploring the vast ternary space, a comprehensive ternary phase diagram remains elusive .
Herein, we conduct a variable stoichiometry structure search at 300 GPa. For the composition space, we restrict the number of Ca and Mg atoms to fewer than 4 each, with the maximum hydrogen-to-metal atomic ratio capped at 16. This resulting in a total of 827 irreducible compositions. Additionally, a constraint of a maximum of 80 atoms per unit cell was imposed, primarily to ensure the inclusion of complex cage-type hydride structures. These settings ensure a sufficiently broad search space, guaranteeing the target structure falls within the search range. Finally, leveraging the high computational efficiency of the ACNN, a total of 5,964,338 structures were optimized during the CSP stage within approximately 2.7 days on a 192-core AMD EPYC 9654 CPU node, achieving a throughput of 474 structures per core-hour.
Following the active learning protocol, DFT SCF calculations were performed on an initial set of 500 randomly generated structures to initiate the computational workflow. The workflow was iteratively executed for a total of 19 cycles. In each iteration, 300 compositions exhibiting the lowest predicted values () were first selected. From each of these compositions, the 3 lowest energy structures () were then chosen as PES sampling candidates for subsequently DFT calculations. In other words, this regulates the computationally expensive yet necessary DFT calculation to a maximum of 900 optimal structures per iteration. The prediction errors for energy, force, and stress, evaluated by comparing the dataset from the -th iteration with the ACNN model trained in the -th iteration, are shown in FIG. 3. This validation scheme quantitatively reflects the performance of the ACNN model in predicting potentially unseen local minima structures. The overall decreasing trend in validation error indicates the model’s improving generalization capability with each iteration. By the final iteration, the ACNN achieved a root mean square error (RMSE) of 44 meV/atom for energy, 283 meV/Å for force, and 14 GPa for cell stresses predictions. This continuous improvement is further evidenced by the progressive increase in the number of uninterrupted structure optimization sequences (as shown in FIG 3(IV)).
Finally, 17, 401 structures were automatically sampled for DFT calculations, including 1, 115 structures necessary to correct unphysical descriptions of the ACNN model, and the remaining 16, 286 nearly local minima structures identified as high value candidates. Among these, 8, 994 structures exhibited energy above the convex hull () below 100 meV/atom. For convex hull refinement, we performed DFT structural relaxation on the 1,000 most stable configurations, with a maximum of approximately 49 meV/atom. As a result, we obtain the phase diagram shown in FIG 4 (a), where a thermodynamically stable Mg2Ca6H46 is successfully located. This structure features a hydrogen network adopting the ‘MEP’ zeolite-type topology [73], with Mg and Ca occupying the 2a and 6d Wyckoff sites, respectively. Alternatively, it can also be viewed as a Mg-substituted variant of the recently reported Ca8H46 phase[74], suggesting its a potential high-temperature superconductor. Here, we found = 137 K at 300 GPa.
Due to the characteristic features of hydride structures, ideal metal superhydride crystal phases are often accompanied by numerous hydrogen-deficient structures, which typically exhibit comparable formation enthalpies. To efficiently expand the pool of candidate clathrate hydrides, we adopt the screening strategy developed by Jiang et al. [75]. This strategy employs an automated topological analysis to identify and classify prototype hydrogen cage motifs. Interestingly, the result accommodates multiple metastable structural prototypes previously reported in zeolite database[76, 77] or other metal hydride systems[78, 79] (see FIG 4 (b)(c)), marked by magenta hollow diamonds in FIG4(a). Nevertheless, we report a new metastable Mg3Ca4H60 phase with an of 66 meV/atom at 300 GPa. In this structure, Mg and Ca atoms occupy the 6b and 8c Wyckoff sites respectively, forming two unusual Mg-centered cage with point group symmetry and Ca-centered cage with symmetry. The H–H bond lengths within these cages measure 0.94, 1.10, 1.11, 1.17 and 1.19 Å at 300 GPa, comparable to those observed in known superconducting hydrogen-rich clathrates, such as CaH6 (1.24 Å at 150GPa)[37] and LaH10 (1.07 and 1.16 Å at 300 GPa)[68, 34].
Building upon the identified candidates structures, we further evaluated their using the SISSO-based prediction model developed by Jiang et al. [75]. This model leveraging the projected density of states of hydrogen at the Fermi level () as an electronic descriptor, in combination with geometric features of the structure facilitates rapid estimation. Benchmarking indicates a mean absolute error of approximately 38 K. The estimated values for these metastable hydrides are presented in FIG5, highlighting the presence of numerous potential high-temperature superconductors that may be experimentally accessible. These results underscore the effectiveness of our algorithm in identifying complex superconducting hydrides with promising properties.
III.2 Be-P-N-O system
In the second test, we apply our method to a more challenging crystal structure prediction (CSP) task by focusing on quaternary compounds. Quaternary system typically present a significantly more complex structural search space due to their greater compositional and configurational degrees of freedom. A particularly promising application lies in the discovery of nonlinear optical (NLO) crystals suitable for deep ultraviolet (DUV) regions. These crystals generate deep ultraviolet coherent light through frequency conversion, enabling essential applications in ultra high resolution lithography, photochemical synthesis, and high precision micromachining. To date, KBe2BO3F2 (KBBF)[15, 80] remains the only practical DUV NLO crystal. However, its strong layering habit[81] has significantly hindered its further commercialization. In contrast, phosphates exhibit diverse non-centrosymmetric (NCS) structures, short UV cutoff edges, and favorable single-crystal growth behavior, making them as strong candidates for the discovery of new NLO materials[82, 83, 84]. Given these advantages, the Be-P-N-O system is deemed a suitable test case for evaluating our approach.
Following the same procedure used for the above ternary system, we first established the compositional search space. A key distinction, however, is that rigid charge balance rules under ambient pressure can effectively constrain the compositional search space. To ensure both comprehensive and efficient exploration while incorporating experimentally feasible synthesis pathways, we first systematically constructed pseudo-binary phase diagrams based on known binary and ternary compounds, from which all appearing compositions were incorporated into the CSP search space. To further enrich this space, we added all charge neutral compositions with fewer than 32 atoms per unit cell. In total, 373 irreducible compositions were identified and subsequently scaled to multiple formula units, with unit cell sizes constrained to fewer than 60 atoms.
A total of 9,381,987 structures were fully relaxed during the CSP stage by ACNN model, executed over 29 active learning iterations to efficiently explore the complex configurational landscape. The entire relaxation process took approximately 1.8 days on a 192-core AMD 9654 CPU, achieving a speed of about 1131 structures per core-hour. Improved relaxation efficiency relative to the previous ternary system primarily attributed to the reduced average number of atoms per structure, as well as the ambient pressure conditions and larger atomic radii (relative to hydrogen) decrease the number of neighboring atoms, collectively lowering the computational workload. For structure sampling, a more flexible strategy was employed. During the initial five iterations, two distinct structures with the lowest energy () were selected from each of the 270 compositions () with the lowest predicted , aiming to facilitate broader exploration of the compositional space. In subsequent iterations, the sampling parameters were adjusted to and to more densely probe novel structures within the compositional region that had been largely validated as near the convex hull and potentially thermodynamical stable.
The corresponding prediction errors during each iterations are presented in FIG6(a). The quaternary system exhibits a more gradual decline in validation error, consistent with the observed trend reported by Wang et al. [42] when transitioning from binary to ternary systems. This slower convergence likely arises from the increased structural complexity of higher-order systems, which imposes greater demands on both CSP and the generalization capacity of the MLP. The ACNN model ultimately achieves a RMSE of 62 meV/atom for energy, 325 meV/Å for forces, and 2.2 GPa for cell stresses predictions.
A total of 13,214 structures were selected for SCF calculations, comprising 11,460 local minima and 1,754 unphysical configurations mistakenly predicted as low-enthalpy by ACNN. Among these, top 1,000 structures with the lowest values were subjected to full DFT structural optimization to construct a more accurate convex hull surface. The complete quaternary phase diagram is presented in the Supplementary Information. Three thermodynamically stable phases were identified: -BePNO2, -BeP3NO7, and -BeP3N5O. These phases are distributed within the pseudo-ternary compositional space spanned by BeO-P2O5-P3N5, as shown in FIG 6(b). As illustrated in FIG 6(c), the crystal structures of all three phases are composed of P-centered and Be-centered tetrahedra, interconnected through corner sharing to form robust three-dimensional (3D) frameworks with distinctive cavity geometries. This structure structural motif resembles those reported in prior studies[85, 86], where nitrogen atoms partially substitute oxygen within phosphate frameworks, forming PNxO4-x (x = 1, 2, 3, 4) mixed anionic tetrahedra. Such hybrid coordination environments contribute to the structural diversity and potential functionality of these materials.
The electronic band structures of the thermodynamically stable phases are presented in FIG.7, revealing band gaps of 5.85 eV (212 nm), 5.56 (223 nm), and 3.3 (376 nm), as calculated at the PBE level. While it is well-established that PBE tends to underestimate band gaps, the first two phases remain within the DUV spectral range, suggesting their potential for DUV optical applications. Additionally, the resulting structure pool include nearly 3,000 structures with formation enthalpies within , commonly associated with potential experimental synthesizability. A more comprehensive analysis of the physical properties of these metastable candidates will provided in a separate forthcoming work.
IV Conclusions
In summary, we present ACNN integrated with the CSP technique to establish an efficient and robust active learning workflow aimed at addressing the high computational complexity inherent in predicting complex multicomponent systems. The robust active learning scheme and the effectiveness of the ACNN leverages the ability to explore millions of candidate structures across various stoichiometries and large atomic configurations, enable a more comprehensive investigation of complex systems. We first apply the algorithm to the high pressure Mg–Ca–H system as a test case, where we discover a stable phase along with a series of metastable candidates as promising high- superconductors. We further demonstrate the effectiveness of our method on the more complex quaternary Be–P–N–O system. In this case, we identify two potential DUV NLO crystals: -BePNO2 and -BeP3NO7. Both cases not only corroborate the rich diversity of promising compounds in complex multicomponent systems, but also highlight the capability of our algorithm to efficiently discover functional materials within vast and high-dimensional chemical spaces.
The modular design of the active learning framework offers significant flexibility for application and external integration. This enables dynamic adjustments to the sampling strategy and search space during runtime, as demonstrated in the quaternary Be–P–N–O case above. The ACNN software package is released alongside the first-principles simulation code ARES, forming a technically integrated and robust software ecosystem. By facilitating software–hardware co-design and optimization, this integration fully unleashes the potential of ACNN for materials design and downstream property simulations. Such software design principles provide a computational foundation for automated materials discovery.
Acknowledgements.
This work is supported by the National Key R&D Program of China (Grant No. 2023YFB3003001), the National Natural Science Foundation of China (Grant No. 12374008 and 12022408), the Interdisciplinary Integration and Innovation Project of JLU, and Fundamental Research Funds for the Central Universities and the Program for JLU Science and Technology Innovative Research Team (JLUSTIRT). We also thank the Changchun Computing Center and the Eco-Innovation Center for providing comprehensive computing resources and technical support throughout the completion of this work.References
- Green et al. [2014] M. A. Green, A. Ho-Baillie, and H. J. Snaith, The emergence of perovskite solar cells, Nature photonics 8, 506 (2014).
- Ceder et al. [1998] G. Ceder, Y.-M. Chiang, D. Sadoway, M. Aydinol, Y.-I. Jang, and B. Huang, Identification of cathode materials for lithium batteries guided by first-principles calculations, Nature 392, 694 (1998).
- Liu et al. [2020] C. Liu, H. Chen, S. Wang, Q. Liu, Y.-G. Jiang, D. W. Zhang, M. Liu, and P. Zhou, Two-dimensional materials for next-generation computing technologies, Nature Nanotechnology 15, 545 (2020).
- Nørskov et al. [2009] J. K. Nørskov, T. Bligaard, J. Rossmeisl, and C. H. Christensen, Towards the computational design of solid catalysts, Nature chemistry 1, 37 (2009).
- De Leon et al. [2021] N. P. De Leon, K. M. Itoh, D. Kim, K. K. Mehta, T. E. Northup, H. Paik, B. Palmer, N. Samarth, S. Sangtawesin, and D. W. Steuerman, Materials challenges and opportunities for quantum computing hardware, Science 372, eabb2823 (2021).
- Walsh [2015] A. Walsh, The quest for new functionality, Nature chemistry 7, 274 (2015).
- Lu et al. [2021] H. Lu, J. Tournet, K. Dastafkan, Y. Liu, Y. H. Ng, S. K. Karuturi, C. Zhao, and Z. Yin, Noble-metal-free multicomponent nanointegration for sustainable energy conversion, Chemical Reviews 121, 10271 (2021).
- Bednorz and Müller [1986] J. G. Bednorz and K. A. Müller, Possible high t c superconductivity in the ba- la- cu- o system, Zeitschrift für Physik B Condensed Matter 64, 189 (1986).
- Kamihara et al. [2008] Y. Kamihara, T. Watanabe, M. Hirano, and H. Hosono, Iron-based layered superconductor la [o1-x f x] feas (x= 0.05- 0.12) with t c= 26 k, Journal of the American Chemical Society 130, 3296 (2008).
- Sun et al. [2023] H. Sun, M. Huo, X. Hu, J. Li, Z. Liu, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, et al., Signatures of superconductivity near 80 k in a nickelate under high pressure, Nature 621, 493 (2023).
- Song et al. [2023] Y. Song, J. Bi, Y. Nakamoto, K. Shimizu, H. Liu, B. Zou, G. Liu, H. Wang, and Y. Ma, Stoichiometric ternary superhydride labeh 8 as a new template for high-temperature superconductivity at 110 k under 80 gpa, Physical Review Letters 130, 266001 (2023).
- Hart et al. [2021] G. L. Hart, T. Mueller, C. Toher, and S. Curtarolo, Machine learning for alloys, Nature Reviews Materials 6, 730 (2021).
- Tsai and Yeh [2014] M.-H. Tsai and J.-W. Yeh, High-entropy alloys: a critical review, Materials Research Letters 2, 107 (2014).
- Mo et al. [2012] Y. Mo, S. P. Ong, and G. Ceder, First principles study of the li10gep2s12 lithium super ionic conductor material, Chemistry of Materials 24, 15 (2012).
- Mei et al. [1993] L. Mei, Y. Wang, C. Chen, and B. Wu, Nonlinear optical materials based on mbe2bo3f2 (m= na, k), Journal of applied physics 74, 7014 (1993).
- Chen et al. [1989] C. Chen, Y. Wu, A. Jiang, B. Wu, G. You, R. Li, and S. Lin, New nonlinear-optical crystal: Lib3o5, Journal of the Optical Society of America B 6, 616 (1989).
- Nikogosyan [1991] D. Nikogosyan, Beta barium borate (bbo) a review of its properties and applications, Applied Physics A 52, 359 (1991).
- Xie et al. [2023] C. Xie, E. Tikhonov, D. Chu, M. Wu, I. Kruglov, S. Pan, and Z. Yang, A prediction-driven database to enable rapid discovery of nonlinear optical materials, Science China Materials 66, 4473 (2023).
- Oganov and Valle [2009] A. R. Oganov and M. Valle, How to quantify energy landscapes of solids, The Journal of chemical physics 130 (2009).
- Villars and Iwata [2013] P. Villars and S. Iwata, Pauling file verifies/reveals 12 principles in materials science supporting four cornerstones given by nature, Chemistry of metals and alloys , 81 (2013).
- Butler et al. [2016] K. T. Butler, J. M. Frost, J. M. Skelton, K. L. Svane, and A. Walsh, Computational materials design of crystalline solids, Chemical Society Reviews 45, 6138 (2016).
- Pickard and Needs [2006] C. J. Pickard and R. Needs, High-pressure phases of silane, Physical review letters 97, 045504 (2006).
- Pickard and Needs [2011] C. J. Pickard and R. J. Needs, Ab initio random structure searching, Journal of Physics: Condensed Matter 23, 053201 (2011).
- Oganov and Glass [2006] A. R. Oganov and C. W. Glass, Crystal structure prediction using ab initio evolutionary techniques: Principles and applications, The Journal of chemical physics 124 (2006).
- Oganov et al. [2019] A. R. Oganov, C. J. Pickard, Q. Zhu, and R. J. Needs, Structure prediction drives materials discovery, Nature Reviews Materials 4, 331 (2019).
- Wang et al. [2010] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Crystal structure prediction via particle-swarm optimization, Physical Review B—Condensed Matter and Materials Physics 82, 094116 (2010).
- Wang et al. [2012a] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Calypso: A method for crystal structure prediction, Computer Physics Communications 183, 2063 (2012a).
- Shi et al. [2018a] X. Shi, C. He, C. J. Pickard, C. Tang, and J. Zhong, Stochastic generation of complex crystal structures combining group and graph theory with application to carbon, Phys. Rev. B 97, 014104 (2018a).
- Wang et al. [2023] J. Wang, H. Gao, Y. Han, C. Ding, S. Pan, Y. Wang, Q. Jia, H.-T. Wang, D. Xing, and J. Sun, Magus: machine learning and graph theory assisted universal structure searcher, National Science Review 10, nwad128 (2023).
- Zhu et al. [2022] Q. Zhu, B. Kang, and K. Parrish, Symmetry relation database and its application to ferroelectric materials discovery, MRS Communications 12, 686 (2022).
- Fredericks et al. [2021] S. Fredericks, K. Parrish, D. Sayre, and Q. Zhu, Pyxtal: A python library for crystal structure generation and symmetry analysis, Computer Physics Communications 261, 107810 (2021).
- Shi et al. [2018b] X. Shi, C. He, C. J. Pickard, C. Tang, and J. Zhong, Stochastic generation of complex crystal structures combining group and graph theory with application to carbon, Physical Review B 97, 014104 (2018b).
- Ma et al. [2009] Y. Ma, M. Eremets, A. R. Oganov, Y. Xie, I. Trojan, S. Medvedev, A. O. Lyakhov, M. Valle, and V. Prakapenka, Transparent dense sodium, Nature 458, 182 (2009).
- Liu et al. [2017] H. Liu, I. I. Naumov, R. Hoffmann, N. Ashcroft, and R. J. Hemley, Potential high-tc superconducting lanthanum and yttrium hydrides at high pressure, Proceedings of the National Academy of Sciences 114, 6990 (2017).
- Drozdov et al. [2019] A. Drozdov, P. Kong, V. Minkov, S. Besedin, M. Kuzovnikov, S. Mozaffari, L. Balicas, F. F. Balakirev, D. Graf, V. Prakapenka, et al., Superconductivity at 250 k in lanthanum hydride under high pressures, Nature 569, 528 (2019).
- Somayazulu et al. [2019] M. Somayazulu, M. Ahart, A. K. Mishra, Z. M. Geballe, M. Baldini, Y. Meng, V. V. Struzhkin, and R. J. Hemley, Evidence for superconductivity above 260 k in lanthanum superhydride at megabar pressures, Physical review letters 122, 027001 (2019).
- Wang et al. [2012b] H. Wang, J. S. Tse, K. Tanaka, T. Iitaka, and Y. Ma, Superconductive sodalite-like clathrate calcium hydride at high pressures, Proceedings of the National Academy of Sciences 109, 6463 (2012b).
- Ma et al. [2022] L. Ma, K. Wang, Y. Xie, X. Yang, Y. Wang, M. Zhou, H. Liu, X. Yu, Y. Zhao, H. Wang, et al., High-temperature superconducting phase in clathrate calcium hydride cah 6 up to 215 k at a pressure of 172 gpa, Physical Review Letters 128, 167001 (2022).
- Li et al. [2022] Z. Li, X. He, C. Zhang, X. Wang, S. Zhang, Y. Jia, S. Feng, K. Lu, J. Zhao, J. Zhang, et al., Superconductivity above 200 k discovered in superhydrides of calcium, Nature communications 13, 2863 (2022).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Physical review letters 77, 3865 (1996).
- Deringer et al. [2018] V. L. Deringer, C. J. Pickard, and G. Csányi, Data-driven learning of total and local energies in elemental boron, Physical review letters 120, 156001 (2018).
- Wang et al. [2024] Z. Wang, X. Wang, X. Luo, P. Gao, Y. Sun, J. Lv, H. Wang, Y. Wang, and Y. Ma, Concurrent learning scheme for crystal structure prediction, Physical Review B 109, 094117 (2024).
- Podryabinkin et al. [2019] E. V. Podryabinkin, E. V. Tikhonov, A. V. Shapeev, and A. R. Oganov, Accelerating crystal structure prediction by machine-learning interatomic potentials with active learning, Physical Review B 99, 064114 (2019).
- Merchant et al. [2023] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80 (2023).
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, et al., A foundation model for atomistic materials chemistry, arXiv preprint arXiv:2401.00096 (2023).
- Yang et al. [2024] H. Yang, C. Hu, Y. Zhou, X. Liu, Y. Shi, J. Li, G. Li, Z. Chen, S. Chen, C. Zeni, et al., Mattersim: A deep learning atomistic model across elements, temperatures and pressures, arXiv preprint arXiv:2405.04967 (2024).
- Xie et al. [2024] F. Xie, T. Lu, S. Meng, and M. Liu, Gptff: A high-accuracy out-of-the-box universal ai force field for arbitrary inorganic materials, Science Bulletin 69, 3525 (2024).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
- Deng et al. [2023] B. Deng, P. Zhong, K. Jun, J. Riebesell, K. Han, C. J. Bartel, and G. Ceder, Chgnet as a pretrained universal neural network potential for charge-informed atomistic modelling, Nature Machine Intelligence 5, 1031 (2023).
- Zhang et al. [2024] D. Zhang, X. Liu, X. Zhang, C. Zhang, C. Cai, H. Bi, Y. Du, X. Qin, A. Peng, J. Huang, et al., Dpa-2: a large atomic model as a multi-task learner, npj Computational Materials 10, 293 (2024).
- Wang et al. [2025] X. Wang, C. Zhang, Z. Wang, H. Liu, J. Lv, H. Wang, Y. Ma, et al., Discovery of high-temperature superconducting ternary hydrides via deep learning, arXiv preprint arXiv:2502.16558 (2025).
- Behler and Parrinello [2007] J. Behler and M. Parrinello, Generalized neural-network representation of high-dimensional potential-energy surfaces, Physical review letters 98, 146401 (2007).
- Behler [2011] J. Behler, Atom-centered symmetry functions for constructing high-dimensional neural network potentials, The Journal of chemical physics 134 (2011).
- Bartók et al. [2010] A. P. Bartók, M. C. Payne, R. Kondor, and G. Csányi, Gaussian approximation potentials: The accuracy of quantum mechanics, without the electrons, Physical review letters 104, 136403 (2010).
- Bartók et al. [2013] A. P. Bartók, R. Kondor, and G. Csányi, On representing chemical environments, Physical Review B—Condensed Matter and Materials Physics 87, 184115 (2013).
- Shapeev [2016] A. V. Shapeev, Moment tensor potentials: A class of systematically improvable interatomic potentials, Multiscale Modeling & Simulation 14, 1153 (2016).
- Zhang et al. [2018] L. Zhang, J. Han, H. Wang, W. Saidi, R. Car, et al., End-to-end symmetry preserving inter-atomic potential energy model for finite and extended systems, Advances in neural information processing systems 31 (2018).
- Fan et al. [2022] Z. Fan, Y. Wang, P. Ying, K. Song, J. Wang, Y. Wang, Z. Zeng, K. Xu, E. Lindgren, J. M. Rahm, et al., Gpumd: A package for constructing accurate machine-learned potentials and performing highly efficient atomistic simulations, The Journal of Chemical Physics 157 (2022).
- Batzner et al. [2022] S. Batzner, A. Musaelian, L. Sun, M. Geiger, J. P. Mailoa, M. Kornbluth, N. Molinari, T. E. Smidt, and B. Kozinsky, E (3)-equivariant graph neural networks for data-efficient and accurate interatomic potentials, Nature communications 13, 2453 (2022).
- Drautz [2019] R. Drautz, Atomic cluster expansion for accurate and transferable interatomic potentials, Physical Review B 99, 014104 (2019).
- Vaswani [2017] A. Vaswani, Attention is all you need, Advances in Neural Information Processing Systems (2017).
- Gilmer et al. [2017] J. Gilmer, S. S. Schoenholz, P. F. Riley, O. Vinyals, and G. E. Dahl, Neural message passing for quantum chemistry, in International conference on machine learning (PMLR, 2017) pp. 1263–1272.
- Loshchilov and Hutter [2017] I. Loshchilov and F. Hutter, Decoupled weight decay regularization, arXiv preprint arXiv:1711.05101 (2017).
- Röcken and Zavadlav [2024] S. Röcken and J. Zavadlav, Accurate machine learning force fields via experimental and simulation data fusion, npj Computational Materials 10, 69 (2024).
- Xu et al. [2019] Q. Xu, S. Wang, L. Xue, X. Shao, P. Gao, J. Lv, Y. Wang, and Y. Ma, Ab initio electronic structure calculations using a real-space chebyshev-filtered subspace iteration method, Journal of Physics: Condensed Matter 31, 455901 (2019).
- Hamann [2013] D. Hamann, Optimized norm-conserving vanderbilt pseudopotentials, Physical Review B—Condensed Matter and Materials Physics 88, 085117 (2013).
- Van Setten et al. [2018] M. J. Van Setten, M. Giantomassi, E. Bousquet, M. J. Verstraete, D. R. Hamann, X. Gonze, and G.-M. Rignanese, The pseudodojo: Training and grading a 85 element optimized norm-conserving pseudopotential table, Computer Physics Communications 226, 39 (2018).
- Peng et al. [2017] F. Peng, Y. Sun, C. J. Pickard, R. J. Needs, Q. Wu, and Y. Ma, Hydrogen clathrate structures in rare earth hydrides at high pressures: possible route to room-temperature superconductivity, Physical review letters 119, 107001 (2017).
- Hong et al. [2020] F. Hong, L. Yang, P. Shan, P. Yang, Z. Liu, J. Sun, Y. Yin, X. Yu, J. Cheng, and Z. Zhao, Superconductivity of lanthanum superhydride investigated using the standard four-probe configuration under high pressures, Chinese physics letters 37, 107401 (2020).
- Snider et al. [2021] E. Snider, N. Dasenbrock-Gammon, R. McBride, X. Wang, N. Meyers, K. V. Lawler, E. Zurek, A. Salamat, and R. P. Dias, Synthesis of yttrium superhydride superconductor with a transition temperature up to 262 k by catalytic hydrogenation at high pressures, Physical review letters 126, 117003 (2021).
- Sukmas et al. [2020] W. Sukmas, P. Tsuppayakorn-aek, U. Pinsook, and T. Bovornratanaraks, Near-room-temperature superconductivity of mg/ca substituted metal hexahydride under pressure, Journal of Alloys and Compounds 849, 156434 (2020).
- Cai et al. [2023] W. Cai, V. S. Minkov, Y. Sun, P. Kong, K. Sawchuk, B. Maiorov, F. F. Balakirev, S. Chariton, V. B. Prakapenka, Y. Ma, et al., Superconductivity above 180 k in ca-mg ternary superhydrides at megabar pressures, arXiv preprint arXiv:2312.06090 (2023).
- Baerlocher and McCusker [2017a] C. Baerlocher and L. B. McCusker, Database of zeolite structures (2017a), accessed: March 25, 2025.
- An et al. [2024] D. An, D. Duan, Z. Zhang, Q. Jiang, T. Ma, Z. Huo, H. Song, and T. Cui, Type-i clathrate calcium hydride and its hydrogen-vacancy structures at high pressure, Physical Review B 110, 054505 (2024).
- Jiang et al. [2025] B. Jiang, X. Luo, Y. Sun, X. Zhong, J. Lv, Y. Xie, Y. Ma, and H. Liu, Data-driven search for high-temperature superconductors in ternary hydrides under pressure, Physical Review B 111, 054505 (2025).
- Baerlocher and McCusker [2017b] C. Baerlocher and L. B. McCusker, Database of zeolite structures, http://www.iza-structure.org/databases/ (2017b).
- [77] Reticular Chemistry Structure Resource, Reticular chemistry structure resource, http://rcsr.anu.edu.au/.
- Sun et al. [2019] Y. Sun, J. Lv, Y. Xie, H. Liu, and Y. Ma, Route to a superconducting phase above room temperature in electron-doped hydride compounds under high pressure, Physical review letters 123, 097001 (2019).
- He et al. [2024] X.-L. He, W. Zhao, Y. Xie, A. Hermann, R. J. Hemley, H. Liu, and Y. Ma, Predicted hot superconductivity in lasc2h24 under pressure, Proceedings of the National Academy of Sciences 121, e2401840121 (2024).
- Chen et al. [2009] C. Chen, G. Wang, X. Wang, and Z. Xu, Deep-uv nonlinear optical crystal kbe 2 bo 3 f 2—discovery, growth, optical properties and applications, Applied Physics B 97, 9 (2009).
- Peng et al. [2018] G. Peng, N. Ye, Z. Lin, L. Kang, S. Pan, M. Zhang, C. Lin, X. Long, M. Luo, Y. Chen, et al., Nh4be2bo3f2 and -be2bo3f: Overcoming the layering habit in kbe2bo3f2 for the next-generation deep-ultraviolet nonlinear optical materials, Angewandte Chemie 130, 9106 (2018).
- Zhao et al. [2014] S. Zhao, P. Gong, S. Luo, L. Bai, Z. Lin, C. Ji, T. Chen, M. Hong, and J. Luo, Deep-ultraviolet transparent phosphates rbba2 (po3) 5 and rb2ba3 (p2o7) 2 show nonlinear optical activity from condensation of [po4] 3–units, Journal of the American Chemical Society 136, 8560 (2014).
- Yu et al. [2017] H. Yu, J. Young, H. Wu, W. Zhang, J. M. Rondinelli, and P. S. Halasyamani, M4mg4 (p2o7) 3 (m= k, rb): structural engineering of pyrophosphates for nonlinear optical applications, Chemistry of Materials 29, 1845 (2017).
- Li et al. [2016] L. Li, Y. Wang, B.-H. Lei, S. Han, Z. Yang, K. R. Poeppelmeier, and S. Pan, A new deep-ultraviolet transparent orthophosphate lics2po4 with large second harmonic generation response, Journal of the American Chemical Society 138, 9101 (2016).
- Jin et al. [2023] W. Jin, C. Xie, X. Hou, M. Cheng, E. Tikhonov, M. Wu, S. Pan, and Z. Yang, From monofluorophosphates a2po3f to difluorophosphates apo2f2 (a= alkali metal): Design of a potential deep-ultraviolet nonlinear optical materials system with a shortened phase-matching wavelength, Chemistry of Materials 35, 5281 (2023).
- Xie et al. [2022] C. Xie, A. Tudi, and A. R. Oganov, Pno: a promising deep-uv nonlinear optical material with the largest second harmonic generation effect, Chemical Communications 58, 12491 (2022).